# Faculty of Computer Science, TU Dresden, Germany

### v2

@@ Abstract
We introduce the Riskman ontology & shapes for representing and analysing information about risk management for medical devices.
Risk management is concerned with taking necessary precautions so a medical device does not cause harms for users or the environment.
To date, risk management documentation is submitted to notified bodies (for certification) in the form of semi-structured natural language text.
We propose to use classes from the Riskman ontology to logically model risk management documentation, and to use the included SHACL constraints to check for syntactic completeness and conformity to relevant standards.
In particular, the ontology is modelled after ISO 14971 and the recently published VDE Spec 90025.
Our proposed methodology has the potential to save many person-hours for both manufacturers (when creating risk management documentation) as well as notified bodies (when assessing submitted applications for certification), and thus offers considerable benefits for healthcare and, by extension, society as a whole.
Resource type: OWL EL Ontology, SHACL Shapes
URL: https://w3id.org/riskman/
@@
Risk management Safety assurance OWL EL SHACL1 Introduction
@@
Medical devices typically are safety-critical, meaning their failure under certain conditions can lead to harm to humans or the environment. To ensure that potential harms are minimised, legislation requires manufacturers of medical devices to provide a comprehensive justification that their product is acceptably safe. In Europe, such legal requirements mainly stem from the European Union’s Medical Device Regulation (EU MDR) [1]. After its introduction in 2017, the original plan was that manufacturers have time until May 2024 to re-certify their devices under the new regulations. This caused an immense backlog for notified bodies (organisations certifying medical devices in Europe, e.g., TÜVs in Germany), so the EU decided to extend the transition period until at least 2027 to avert a potential shortage of medical devices (and ramifications thereof) [44].
@@
The fundamental challenges of medical device certification are (1) the sheer amount of information that notified bodies have to process (even for a single device), and (2) the way this information is submitted to notified bodies, namely in the form of text. In principle, device manufacturers have to argue that they have proactively, systematically, and thoroughly analysed and mitigated the risks associated with their device as much as possible. One way to do this is using a so-called assurance case, a structured argument supported by a body of evidence that together provide a compelling, comprehensible, and valid case that a system is acceptably safe for a given application in a given context. Using assurance cases has a long(er) tradition in other areas of safety-critical systems, e.g. in aviation or nuclear power plants [67], but the need for assurance cases for medical devices has been recognised long ago [73]. The Food and Drug Administration of the U.S. (FDA) has even provided a guidance document for one class of products (infusion pumps) due to an unusually large number of previous incidents [69].
@@
As much as the introduction of assurance cases has helped advance safety management practices in the medical device industry, there is still a lot to be wanted. Sujan et al. even claim that “safety management practices in healthcare [are] at present […] less mature than those in other safety-critical industries” [67, p. 185]. We briefly note some of the shortcomings we address in this work.
@@
Since the intended recipient of a safety case is a human (auditor), a lot of work in safety engineering has focused on the presentation of the argument, e.g. to avoid confirmation bias [17]. In particular, several approaches to visualise assurance cases exist, e.g. the goal structuring notation [46]. There is much less work on the representation of safety arguments, since applicable legislation typically requires a submission in the form of text.
@@
Due to the sheer volume of text submitted to notified bodies for certification, an auditor can never look at all relevant points in detail; furthermore, auditors typically spend a considerable proportion of their time navigating through documents (best case, using text search in a document viewer). Significant time savings are offered by semi-automatic checks and improved navigation through e.g. semantics-enhanced document search.
@@
Reusing parts of safety arguments is a widely recognised problem [52, 60]. The simplest instance of re-use happens when a manufacturer continues development of a certified device and wants to certify the “updated version”: application lifecycle management tools can help identify the delta between two device versions and risk managers only need to concentrate on how that difference affects risks and mitigations. Harder instances of re-use are equally pervasive and important, e.g. when a manufacturer wants to use a part of a device in a novel device of a different type, or when a manufacturer buys parts (like integrated circuits) from component suppliers and needs to use assurances about the parts in order to claim overall device safety. We even propose to go beyond this and technically enable re-use across manufacturers, which is especially relevant for small/medium enterprises that may lack staff or experience in risk management.
Thus there is substantial demand for intelligent software support in risk management, both on the side of manufacturers as well as from notified bodies.
@@
In this paper, we propose to use logical modelling to represent risk management documentation, and the shapes constraint language (SHACL) to check those representations for conformance with a set of requirements, e.g., whether all identified risks have an associated mitigation. While such straightforward checks could clearly also be achieved by custom software, the presence of a logical inference step before the constraint checking step is a significant advantage of our approach, as we shall demonstrate later in the paper. Taken together, the Riskman ontology and its shapes constraints constitute the resource we contribute in this paper, which is intended to be Riskman’s canonical citation. As far as we are aware, using OWL & SHACL is a novel approach to representing and reasoning about risk management documentation. The requirements to be checked against are simple and syntactic at the moment, but from communication with domain experts from notified bodies and manufacturers, we still expect our contributed resource and methodology to be a major step forward as manufacturers (and notified bodies) can expect to “save” the first round of conformity assessment, which can translate into a several-months-long reduction in time to market. On the other hand, our resource does not aim at assessing the adequacy and correctness of the implemented mitigations (such as whether an insulation thickness of 0.5mm is adequate to lower the risk of electric shock), which is presumably AI-complete and would require vast amounts of background knowledge. Apart from this important application domain, our ontology and shapes are also of independent interest to the Semantic Web community, as the combination of logical inference and SHACL constraint checking is a topic that has garnered research interest lately [4, 57, 22]. Finally, the use of our approach in the medical device industry and on the part of notified bodies entails a considerable impact on the adoption of Semantic Web technologies in general. By employing these technologies ourselves, we furthermore enable Riskman users to benefit from the good infrastructure of tooling and data the Semantic Web field provides. Risk reports can easily be queried or enhanced with additional information about the use context of a concrete device or data about particular patient problems by simply using other ontologies and data [55, 25, 54].
@@
The rest of the paper is organised as follows: In the next section, we give an overview on related work, mostly ontologies involving notions of risk, but also software-based solutions for safety case management. Afterwards (Sect. 3), we introduce the state of the art of risk management and the basic notions introduced by the norms upon which we build our work, both in general terms and with an illustrative running example. Sect. 4 then introduces the Riskman ontology and its associated shape constraints, explaining how we intend them to be used in risk management, and showcasing them with the running example. In Sect. 5 we conclude with a discussion of potential future work.
@@ 2.1 Conceptual Work and Ontologies
Fenz et al. [27, 26] created an ontology for the information security standard ISO 27002. While they do not deal with risk management (or medical devices), they also use their ontology along with a reasoner to infer information about compliance with the standard’s requirements. Reasoning results are then interpreted by humans or by a tailor-made software tool.
Uciteli et al. [68] provided the Risk Identification Ontology, which defines notions of risk and adverse situation and embeds them into the top-level ontology GFO (General Formal Ontology) [37, 38]. However, they work in a more process-oriented setting with special focus on risk identification in time periods surrounding surgical procedures. In that setting, risks cannot be mitigated beforehand, and so risk management – our main focus – is not within their scope.
Kim et al. [47] presented a process integration ontology for medical software developers with a focus on medical devices, combining notions from IEC 60601-1, IEC 62304, and, notably, ISO 14971. The resulting integrated ontology is however mainly designed to help developers comply with the involved standards; the ontology was not developed with logical reasoning as explicit intended use case.
Aziz et al. [9] developed a Hazard Identification Ontology, involving notions of Hazard and Events. Their focus was however on identifying risks rather than mitigating them, especially in scenarios of fire, explosions, or toxicity.
Schütz et al. [62] created an ontology for medical devices in Germany, albeit more broadly targeting devices’ manufacturers, operators, and legal procedures from an outside perspective with the aim of general semantic interoperability, and based on a legal framework that has since been superseded by the EU MDR.
Single et al. [64] presented an ontology for Hazard and Operability (HAZOP), a methodology for scenario-based hazard evaluation, therein defining notions of deviation, cause, effect, consequence, and safeguard. Their aim was however to create HAZOP worksheets (to be used by human operators) automatically.
Alanen et al. [5] provided a comprehensive risk assessment ontology including the notions risk, risk level, and risk control, which is harmonised between (and intended for use across) safety, security, and dependability. The intended use of their ontology is to “support the creation of a structured work product storage with traceability links” in order to improve upon current practices with non-structured word-processing documents, albeit without using logical reasoning. Conclusion: While various ontologies including notions of risk exist, they either do not implement ISO 14971 [40] or are not designed with logical reasoning in mind, and furthermore none of them uses SHACL for conformance testing.
@@
Fujita et al. [29] presented the D-Case Editor, a dependability-oriented assurance case editor supporting the goal structuring notation [46], implemented as an Eclipse plugin and available at its GitHub repository [20]. However, its development has been discontinued (the last repository change was in 2015).
Rushby [61] and Cruanes et al. [19] introduced the Evidential Tool Bus, a Datalog-based system for integrating the development of safety cases into system (and software) development. In particular, their approach covers the management of claims (about the system to be developed) and how they are supported by evidence provided by other software tools. While an implementation is available [56], it is not maintained any longer (last change 2016). Beyene and Ruess [13] later picked up that work and extended it by connecting to the Jenkins [42] continuous integration software, with an active repository [12]. The focus of the Evidential Tool Bus is primarily on obtaining, maintaining, and managing pieces of evidence in a development setting. As such, it can be an important and complementary addition to using the methodology we propose in this paper.
Since the issue of safety is of great concern to industry, there are also several proprietary software tools centring around safety assurance: The UK company Adelard offers the Assurance and Safety Case Environment. Similarly, de la Vara et al. [21] reported on using the V&V studio by REUSE Software.
Conclusion: Existing software solutions for risk management are proprietary or do not use formal logics, or they have other (but complementary) use cases.
@@ 3 Background
| Analysed risk | Combination of one or more domain-specific hazard(s) with one hazardous situation and one harm with reference to a device context and a specification of an initial risk level. |
| Controlled risk | Structured artifact that relates one analysed risk with one or more SDA(s) and specifies a resulting residual risk. |
@@
The Riskman ontology formalises the recent VDE Spec 90025 [7], which proposes a structured format for digitalising risk management files, as well as a machine-readable exchange format using HTML and RDFa [2]. VDE Spec 90025, in turn, is based on ISO 14971 [40], which “specifies terminology, principles and a process for risk management of medical devices”. In this section, we briefly explain the notions from VDE Spec 90025 and ISO 14971 that are central to the Riskman ontology. Some of the notions are provided in Table 1. In the supplementary material we give a full list, as VDE Spec 90025 [7] is still in press.
The model of risk underlying both ISO 14971 and VDE Spec 90025 (see Annex C of ISO 14971) is that of a hazard leading, via a sequence of events, to a hazardous situation, which, in turn, results in a harm, as depicted below:
The risk analysis that manufacturers of medical devices must undertake involves, in this view, compiling a list of known and foreseeable combinations of hazard, events, hazardous situation, and harm. A further central aspect of risk assessment is estimating (initial) risk levels by associating a probability of occurrence and severity to harms. The probability is often split into probabilities P1 and P2, with the overall probability then being .
Building on ISO 14971, VDE Spec 90025 requires manufacturers to document the results of their risk analysis as a list of analysed risks. This includes providing a device context to each analysed risk. Moreover, hazards are embedded within domain-specific hazards, which also include information on the device function and device component of the device (or class of devices from one domain, e.g. X-ray devices) to which the hazard refers. VDE Spec 90025 does not provide requirements on the values used to specify probabilities or the severity; in particular, probabilities are often not probabilities in the mathematical sense. Rather, these values are interpreted as magnitudes within a scale that is used by the manufacturing company producing the risk management file (cf. Sect. 4.4).
@@
Risk assessment by manufacturers primarily aims at risk control, i.e. minimising risks as much as possible. To document the mitigation strategies devised for the different risks, VDE Spec 90025 proposes the use of safe design arguments (SDAs). These, although building on the notion of assurance cases [73], avoid some of the complexities in their use [45, 51, 31, 32] by only requiring a considerably simplified structure. Specifically, SDAs encode structured arguments, showing that a certain risk has been mitigated, as trees. Thus, each SDA can have one or more sub-SDAs (its children), which serve to substantiate a claim made in the parent SDA. Moreover, SDAs need ultimately be based on safe design argument implementations (SDAIs), i.e. all leaf SDAs need to be SDAIs. These include not only a claim but also an implementation manifest. The latter gives detailed information on how the claim has been implemented and points to concrete evidence (e.g. additional documentation) to support this.
@@
VDE Spec 90025 makes a further distinction between risk SDAs and assurance SDAs, with assurance SDAs referring to some state-of-the-art safety assurance – e.g. a section of a norm or standard mentioning a way of handling a risk. All other SDAs are risk SDAs, while children of assurance SDAs need also be assurance SDAs. An SDA tree is embedded within a controlled risk, which, apart from referring to the analysed risk it “controls”, also indicates a residual risk level, i.e. probability and severity of the risk remaining after risk control.
@@
Insulin infusion pumps aid in regulating blood glucose levels, especially of patients with diabetes, by administering fast-acting insulin via a catheter inserted beneath the skin. Based on the risk assessment for a generic infusion pump by Zhang et al. [75, 74], Fig. 1 shows a controlled risk and associated SDA that can be extracted from a risk management file that follows VDE Spec 90025.
forked edges,
for tree=align=center,base=bottom,draw,font=,edge=-¿,rounded corners,scale=0.9,
im/.style=fill=green!10,
sa/.style=fill=purple!15,
sda/.style=fill=blue!10,
asda/.style=fill=brown!15,
[SDA:
Alternative alerting when vibration mechanism of non-audio alarm fails (), sda,
[SDA:Additional
visual (blinking)
signal (),sda,name=sda1
[IM: Sec. 10.3
of Alarm
report (),im,name=im1]],
[SDA: Notification
recurs every
minutes (),sda,name=sda2
[IM: ,
Sec. 10.7
of Alarm
report (),im]],
[SDA: Additio-
nal audio
alarm (),sda,name=asda3
[SDA: Audio
signal if vibration
signal not
acknowledged (),sda,name=asda4
[IM: Sec.
10.11 of Alarm
report (),im]],
[SDA:
Audible signal is
at least X db
at Y m (),asda,name=asda5
[IM: ,
, Sec. 5.3
of Loudspeaker
test (),im,name=im5]],
[SA: IEC
60601 (), no edge, sa,name=sa5,yshift=6pt]
[…]
4 The Riskman Ontology & Shapes: Overview and Usage

### v3

# Supporting Risk Management for Medical Devices via the Riskman Ontology and Shapes (Preprint)

@@ Abstract
We propose the Riskman ontology and shapes for representing and analysing information about risk management for medical devices. Risk management is concerned with taking necessary precautions to ensure that a medical device does not cause harms for users or the environment. To date, risk management documentation is submitted to notified bodies (for certification) in the form of semi-structured natural language text. We propose to use terms from the Riskman ontology to provide a formal, logical underpinning for risk management documentation, and to use the included SHACL constraints to check whether the provided data is in accordance with the requirements of the two relevant norms, i.e. ISO 14971 and VDE Spec 90025. Riskman is available at https://w3id.org/riskman.
@@
[…]
@@
Medical devices typically are safety-critical, meaning their failure under certain conditions can lead to harm to humans or the environment. To ensure that potential harms are minimized, legislation requires manufacturers of medical devices to provide a comprehensive justification that their product is acceptably safe. In Europe, such legal requirements mainly stem from the European Union’s Medical Device Regulation (EU MDR) [EUMDR]. After its introduction in 2017, the original plan was that manufacturers have time until May 2024 to re-certify their devices under the new regulations. This caused an immense backlog for notified bodies (organisations certifying medical devices in Europe, e.g., TÜVs in Germany), so the EU decided to extend the transition period until at least 2027 to avert a potential shortage of medical devices (and ramifications thereof) [mdrNewsURL].
@@
The fundamental challenges of medical device certification are (1) the sheer amount of information that notified bodies have to process (even for a single device), and (2) the way this information is submitted to notified bodies, namely in the form of text. In principle, device manufacturers have to argue that they have proactively, systematically, and thoroughly analyzed and mitigated the risks associated with their device as much as possible. One way to do this is using a so-called assurance case (AC), a structured argument supported by a body of evidence that jointly provide a compelling, comprehensible, and valid case that a system is acceptably safe for a given application in a given context. Using assurance cases has a long(er) tradition in other areas of safety-critical systems, e.g. in aviation or nuclear power plants [SujanHKPJ16], but the need for assurance cases for medical devices has been recognized long ago [WeinstockG09]. The Food and Drug Administration of the U.S. (FDA) has even provided a guidance document for one class of products (infusion pumps) due to an unusually large number of previous incidents [FDA14].
@@
As much as the introduction of assurance cases has helped advance safety management practices in the medical device industry, there is still a lot to be wanted. Sujan et al. even claim that “safety management practices in healthcare [are] at present […] less mature than those in other safety-critical industries” [SujanHKPJ16, p. 185]. We briefly note some identified shortcomings:
@@
(1) Since the intended recipient of a safety case is human (the auditor), a lot of work in safety engineering has focused on the presentation of the argument, e.g. to avoid confirmation bias [ChowdhuryWPL20]. Several approaches to visualize assurance cases exist, e.g. the goal structuring notation [KellyM97]. However, work on the representation of safety arguments is rare, since applicable legislation typically requires a submission in the form of text.
@@
(2) Due to the sheer volume of text submitted to notified bodies for certification, an auditor can never look at all relevant points in detail; furthermore, auditors typically spend a considerable proportion of their time navigating through documents (in the best case using text search in a document viewer). Tedious, manual examination of submissions is necessary to identify missing or incorrect information. Significant time savings are offered when the initial inspection can be (semi-)automated and when improved navigation through e.g. semantics-enhanced document search is available.
@@
(3) Reusing parts of safety arguments is a widely recognized problem [MartinBLW14, RuizJEVL17]. The simplest instance of re-use happens when a manufacturer continues development of a certified device and wants to certify the “updated version”: application lifecycle management tools help identify the delta between two device versions and risk managers can restrict attention on how that difference affects risks and mitigations. Harder instances of re-use are equally pervasive and important. For example, this occurs when a manufacturer wants to use some device part in a novel device of a different type. It also arises when a manufacturer buys parts (like integrated circuits) from component suppliers and uses assurances about those parts to claim overall device safety. We even propose to go beyond this and technically enable re-use across manufacturers, which is especially relevant for small/medium enterprises that may lack staff or experience in risk management.
Therefore, any approach toward a more structured representation of safety cases, allowing for automated checks and facilitating re-use, would appear to be highly welcome by all involved parties. Consequently, in this paper, we propose to use logical modelling (more precisely the web ontology language, OWL) to represent risk management documentation, and the Shapes Constraint Language (SHACL) to check those representations for conformance with a set of requirements, e.g., whether all identified risks have an associated mitigation. While such straightforward checks could clearly also be achieved by custom software, the presence of a logical inference step before the constraint checking step is a significant advantage of our approach, as we shall demonstrate later in the paper. Thus as our main contribution we introduce the Riskman ontology and shapes.
@@
As far as we are aware, using OWL and SHACL is a novel approach to representing and reasoning about risk management documentation. The requirements to be checked are simple and syntactic at the moment, but in view of communication with domain experts from notified bodies and manufacturers, we still expect our approach to be a major step forward as manufacturers (and notified bodies) can expect to the first round of conformity assessment being fully automated. We would like to emphasize that our approach does not aim at assessing the adequacy and correctness of the implemented mitigations themselves (such as whether an insulation thickness of 0.5 mm is adequate to lower the risk of electric shock). It stands to reason that such an analysis is presumably AI-complete and would at least require vast amounts of background knowledge. Apart from facilitating tasks in this important application domain, our ontology and shapes are also of independent interest to the Semantic Web community, as the combination of logical inference and SHACL constraint checking is a topic that has garnered significant research interest lately [AhmetajOOS23, ParetiKNS19, demeesterswj2021]. Finally, the fact that our approach relies on standardized and well-supported formalisms unlocks the adoption of Semantic Web technologies in general. Employing these technologies enables Riskman users to benefit from existing infrastructure of tooling: Risk reports can easily be queried or enhanced with additional information about the use context of a concrete device or data about particular patient problems by simply using other ontologies and data [cc, el2018snomed, ncitURL].
@@
The rest of the paper is organized as follows: In the next section, we give an overview on related work. Afterwards (Section 3), we introduce the state of the art of risk management and the basic notions introduced by the norms upon which we build our ontology, both in general terms and with an illustrative running example. Section 4 then introduces the Riskman ontology and its associated shape constraints, explaining how we intend them to be used in risk management, and showcasing them with the running example. In Section 5 we conclude with a discussion of potential future work.
@@ 2 Related Work
We reviewed the literature for (1) ontologies on (a) medical devices or (b) notions of risk, with a special focus on an intended use for automated reasoning (in particular using SHACL constraints), and (2) academic reports on software tools for risk management.
@@ 2.1 Conceptual Work and Ontologies
Fenz et al. [FenzPH16, FenzN18] created an ontology for the information security standard ISO 27002. While they do not deal with risk management (or medical devices), they also use their ontology along with a reasoner to infer information about compliance with the standard’s requirements. Reasoning results are then interpreted by humans or by a tailor-made software tool. Uciteli et al. [UciteliNTSSFKSS17] provided the Risk Identification Ontology, which defines notions of risk and adverse situation and embeds them into the top-level ontology GFO (General Formal Ontology) [Herre10, HerreHBHLM06]. However, they work in a more process-oriented setting with special focus on risk identification in time periods surrounding surgical procedures. In that setting, risks cannot be mitigated beforehand, and so risk management – our main focus – is not within their scope. Kim et al. [KimPLL19] presented a process integration ontology for medical software developers with a focus on medical devices, combining notions from IEC 60601-1, IEC 62304, and, notably, ISO 14971. The resulting integrated ontology is however mainly designed to help developers comply with the involved standards; the ontology was not developed with logical reasoning as explicit intended use case. Aziz et al. [AzizAK19] developed a Hazard Identification Ontology, involving notions of Hazard and Events. Their focus was however on identifying risks rather than mitigating them, especially in scenarios of fire, explosions, or toxicity. Schütz et al. [SchutzFW20] created an ontology for medical devices in Germany, albeit more broadly targeting devices’ manufacturers, operators, and legal procedures from an outside perspective with the aim of general semantic interoperability, and based on a legal framework that has since been superseded by the EU MDR. Single et al. [SingleSD20] presented an ontology for Hazard and Operability (HAZOP), a methodology for scenario-based hazard evaluation, therein defining notions of deviation, cause, effect, consequence, and safeguard. Their aim was however to create HAZOP worksheets (to be used by human operators) automatically. McDonald et al. [McDonald21] mention how they use the “ARK Mindful Governance of operational risk formal ontology” to annotate textual risk-analysis data to make it amenable to machine processing (p12). They however give no details on the ontology or its possible inferences. Alanen et al. [AlanenLMPAHT22] provided a comprehensive risk assessment ontology including the notions risk, risk level, and risk control, which is harmonized between (and intended for use across) safety, security, and dependability. The intended use of their ontology is to “support the creation of a structured work product storage with traceability links” in order to improve upon current practices with non-structured word-processing documents, albeit without using logical reasoning. Golpayegani et al. [GolpayeganiPL22] propose the AI Risk Ontology (AIRO) for expressing information associated with high-risk AI systems based on the requirements of the EU’s AI Act [AIAct] and the ISO 31000 series of standards. AIRO contains classes for risk, risk source, event, consequence, and impact, and Golpayegani et al. also present SHACL shapes for determining whether an AI system is high-risk. This is perhaps the closest relative of Riskman, only for AI systems instead of medical devices. In conclusion, while various ontologies including notions of risk exist, they either do not implement ISO 14971 [ISO14971] or are not designed with logical reasoning in mind, and most of them do not use SHACL for conformance testing (in the sense of pre-certification).
@@
Fujita et al. [FujitaMHSKI12] presented the D-Case Editor, a dependability-oriented assurance case editor supporting the goal structuring notation [KellyM97], implemented as an Eclipse plugin and available at its GitHub repository [d-case_editorURL]. However, its development has been discontinued. Rushby [Rushby05] and Cruanes et al. [CruanesHOS13] introduced the Evidential Tool Bus, a Datalog-based system for integrating the development of safety cases into system (and software) development. In particular, their approach covers the management of claims (about the system to be developed) and how they are supported by evidence provided by other software tools. While an implementation is available [ETBURL], it is not maintained any longer. Beyene and Ruess [BeyeneR18] later picked up that work and extended it by connecting to the Jenkins [JenkinsURL] continuous integration software, with an active repository [JenkinsURLGit]. The focus of the Evidential Tool Bus is primarily on obtaining, maintaining, and managing pieces of evidence in a development setting. As such, it can be an important and complementary addition to using the methodology we propose in this paper. Since the issue of safety is of great concern to industry, there are also several proprietary software tools centring around safety assurance: The UK company Adelard offers the Assurance and Safety Case Environment. Similarly, de la Vara et al. [delaVaraJMP19] reported on using the V&V studio by REUSE Software. In conclusion, existing software solutions for risk management are proprietary or do not use formal logics, or they have other (but complementary) use cases.
@@ 3 Background
| Analyzed risk | Combination of one or more domain-specific hazard(s) with one hazardous situation and one harm with reference to a device context and a specification of an initial risk level. |
| Controlled risk | Structured artifact that relates one analyzed risk with one or more SDA(s) and specifies a resulting residual risk. |
@@
The Riskman ontology is based on the recent VDE Spec 90025 [VDESpec24], which proposes a structured format for digitalizing risk management files, as well as a machine-readable exchange format using HTML with RDFa [RDFa], which annotates (some) HTML tags with Resource Description Framework (RDF) triples. VDE Spec 90025, in turn, is based on ISO 14971 [ISO14971], which “specifies terminology, principles, and a process for risk management of medical devices”. Throughout this section, we briefly recall the notions from VDE Spec 90025 and ISO 14971 that are central to the Riskman ontology. Some of the notions are provided in Table 1. A full list of definitions of the terms can be found in VDE Spec 90025 [VDESpec24, Section 3]. The model of risk underlying both ISO 14971 and VDE Spec 90025 (see Annex C of ISO 14971) is that of a hazard leading, via a sequence of events, to a hazardous situation, which, in turn, results in a harm, as depicted below:
According to this view, the risk analysis that manufacturers of medical devices must undertake involves compiling a list of known and foreseeable combinations of hazard, events, hazardous situation, and harm. A further central aspect of risk assessment is estimating (initial) risk levels by associating a probability of occurrence and a severity to each harm. The probability is often split into probabilities and , with the overall probability then being . Building on ISO 14971, VDE Spec 90025 requires manufacturers to document risk analysis results as a list of analyzed risks, including related harm, patient problem, device context, the hazardous situation caused by preceding event(s), the assessed initial risk level, and the associated domain specific hazards. These, in turn, further specify the underlying hazard through details on the device problem, device function, and device component. The “problem” classes enable referencing terminology from the International Medical Device Regulators Forum (IMDRF) [IMDRF20AET], specifically Annex E for patient problems and Annex A for device problems. The notion of a device is absent in VDE Spec 90025 and the Riskman ontology, since each risk management file addresses a single medical device. However, extending the ontology to include a device concept and link related entities would enable more comprehensive queries. VDE Spec 90025 does not provide requirements on the values used to specify probabilities or the severity either; in particular, probabilities are not probabilities in the mathematical sense. Rather, these values are interpreted as magnitudes within a scale that is used by the manufacturing company producing the risk management file (cf. Section 4.3).
@@
Risk assessment by manufacturers primarily aims at risk control, i.e. minimizing risks as much as possible [EUMDR]. To document the mitigation strategies devised for the different risks, VDE Spec 90025 proposes the use of safe design arguments (SDAs). These, although building on the notion of assurance cases [WeinstockG09], avoid some complexity in their use [Kelly03, Leveson11, Graydon15, Graydon17] by only requiring a considerably simplified structure. Specifically, SDAs encode structured arguments, showing that a certain risk has been mitigated, as trees. Thus, each SDA can have one or more sub-SDAs (its children), which serve to substantiate a claim made in the parent SDA. Moreover, SDAs need ultimately be based on safe design argument implementations (SDAIs), i.e. all leaf SDAs need to be SDAIs. These include not only a claim but also an implementation manifest, which gives detailed information on how the claim has been implemented and points to concrete evidence (e.g. additional documentation or a specific line of software source code) to support this.
@@
VDE Spec 90025 further distinguishes between risk SDAs and assurance SDAs, with assurance SDAs referring to some state-of-the-art safety assurance – e.g. a section of a norm mentioning a way of handling a risk. All other SDAs are risk SDAs, while children of assurance SDAs must also be assurance SDAs. An SDA tree is embedded within a controlled risk, which, apart from referring to the analyzed risk it “controls”, also indicates a residual risk level, i.e. probability and severity that remain after risk control.
@@
Insulin infusion pumps aid in regulating blood glucose levels, especially of patients with diabetes, by administering fast-acting insulin via a catheter inserted beneath the skin. Based on the risk assessment for a generic infusion pump by Zhang et al. [zhangJJ10, zhangJJR11], Figure 1 shows a controlled risk and associated SDA that can be extracted from a risk management file that follows VDE Spec 90025. Following Zhang et al. [zhangJJ10, entry 4.3.9 in Table 4 in the appendix], the risk stems from an “alarm malfunction” hazard (indicated by in Figure 1). It is further specified within a domain specific hazard () with related information on the device component (), its function () and problem (). Specifically, the vibration mechanism of the non-audio alarm integrated into the pump may fail (event ) under normal operating conditions (device context ). Then, the patient may not become aware of an issue (event ), which can lead to the patient receiving less insulin (hazardous situation ) and the patient losing consciousness (harm ). The SDA (, based on the work of Zhang et al. [zhangJJR11, Table 3]) consists of three sub-SDAs and expresses that there are alternative means of alerting the patient. The first sub-SDA () expresses that the alarm condition is also indicated through visual signals; the second sub-SDA () indicates that this notification is recurring; the third sub-SDA () expresses that there is also an additional audio alarm that will start unless the patient acknowledges the vibration or blinking. Moreover, according to sub-SDA () the audible signal is in accordance with regulations, here the assurance is IEC 60601 (). Thus, sub-SDA () is the only assurance SDA; all other SDAs are risk SDAs. On the other hand, as required by VDE Spec, each leaf SDA is an SDAI, with their implementation manifests (,,,) pointing to details and documentation.
4 The Riskman Ontology and Shapes: Overview and Usage